# TinySAM: Pushing the Envelope for Efficient Segment Anything Model

### v2

@@ 3.3 Quantization
SAM-H [22] 2976G 2392 46.6 30.8 51.0 61.7 44.7 32.5 57.6 65.5
SAM-L [22] 1491G 1146 46.2 30.2 50.1 60.5 43.5 31.1 56.3 65.1
SAM-B [22] 487G 368.8 43.4 28.5 45.5 53.4 40.8 29.1 52.8 60.7
FastSAM [51] 344G 153.6 37.9 23.9 43.4 50.0 34.5 24.6 46.2 50.8
EfficientSAM-Ti [47] 106G 81.0 42.3 26.7 46.2 57.4 39.9 28.9 51.0 59.9
MobileSAM [50] 42.0G 38.4 41.0 24.4 44.5 58.6 37.0 24.7 47.8 59.1
@@
3.4 Hierarchical Segmenting Everything
SAM proposes an automatic mask generator which samples points as a grid to segment everything. However, we find that dense points grid leads to over-finegrained segmentation results and also occupies massive computing resources. On the one hand, for a complete object, too many sampling points may cause slightly different parts of the object to be incorrectly segmented as separate masks. On the other hand, since the image encoder has been largely shrinked, the time cost of everything mode inference is mainly in the mask decoder part. For the default settings of SAM automatic mask generator, it samples $32\times 32=1024$ points as the prompts, which means the mask decoder is inferred by $1024$ times. It costs $16$ms for image encoder and $894$ms for mask decoder on a single V100 GPU.
To reduce the time cost of everything mode, we propose a hierarchical mask generating method. The comparison between our hierarchical strategy and the original one is shown in Figure 3. Different from original everything mode, in the first step we only use $25\%$ points in each side so the total points is $1/16$ of the original settings, as shown in Figure 3(c). Then we infer the prompt encoder and mask decoder with these prompts and get the results.
We then filter out some masks with confidence exceeding a threshold $\tau$, and mark the corresponding locations as areas that could be considered as final predictions. For these areas, since they are considered as the segmentation results of instances with high confidences, there is no need to re-generate point prompts. Then we sample points as the same density with original settings but ignore points in the above area. As shown in Figure 3(d), most points on the grass and body of the first cow are ignored by our strategy. Meanwhile, the points on the second cow and the sky are kept to further segment. Specifically, the second cow is incorrectly segmented as the same object with the first cow in the initial round. This strategy can avoid both the cost of inference time and the over-finegrained segmentation of the object. Then we utilize the point prompts sampled in the second round to get the mask predictions. Finally, the results of these two round are merged and post-processed to get the final masks. More than $50\%$ points are ignored by our method thus brings in great latency reduction.
4 Experiments
4.1 Implementation Details
We utilize the TinyViT-5M [46] as the lightweight student image encoder and SAM-H as the teacher model, following prior work [50]. Partial SA-1B dataset is used as the training data for full-stage distillation. We adopt Adam optimizer and train the student network for 8 epochs. For each iteration, we sample 64 prompts according to hard prompt sampling strategy described in Section 3.2. To accelerate the distillation process, the image embeddings from the teacher network have been computed and stored in advance. Therefore, the heavy image encoder of teacher network is not necessary to compute repeatedly during training time. For post training quantization, we set $\alpha=0.01,\beta=1.2,n=100,rounds=3$ for iterative search. We calibrate quantized model on SA-1B dataset using 8 images. We conduct zero-shot evaluation on downstream tasks like instance segmentation and point prompt segmentation. Following the suggestions by SAM [21], the multi-output mode is adopted and the final mask prediction is the one with highest IoU prediction.
@@ 4.1 Implementation Details
Figure 4: Evaluation results of zero-shot points valid mask. X-axis represents the number of points used in prompts and Y-axis represents the mIoU across all masks evaluated. The proposed TinySAM outperforms MobileSAM and achieves results close to SAM ViT-B.
4.2 Zero-Shot Instance Segmentation
For zero-shot instance segmentation task, we strictly follow the experimental settings of SAM and use the object detection results of ViTDet-H [23] as the box prompt for instance segmentation. We evaluate the zero-shot instance segmentation task for models on the benchmark of COCO [28] dataset and LVIS v1 [17]. We compare our TinySAM with different variants of SAM [22], and also with prior efficient models like FastSAM [51], MobileSAM [50] and EfficientSAM [47]. As shown in Table 1, the proposed TinySAM obtained superior performance when compared with prior methods. Specifically, our TinySAM outperforms FastSAM [51] in terms of FLOPs and instance segmentation accuracy, i.e., about $4\%$ AP improvement with only $12.2\%$ FLOPs and $25\%$ latency. With the same computational cost, our TinySAM also achieves $1.3\%+$ AP on COCO dataset than MobileSAM [50] and $1.9\%+$ AP on LVIS v1 dataset, respectively. With similar performance on COCO dataset, TinySAM is 2×2$\times$2 × faster than EfficientSAM [47]. Our W8A8 quantized variant of TinySAM (Q-TinySAM) also obtains competitive performance across different methods. Specifically, Q-TinySAM achieves $0.4\%+$ AP on COCO dataset than MobileSAM [50], with only $48\%$ FLOPs and $62.5\%$ latency. Some visual results on COCO validation set are shown in Figure 5. Our proposed TinySAM captures more clear and smooth boundaries compared with other efficient variants of SAM.
4.3 Zero-shot Points Valid Mask Evaluation
@@
In this section we also evaluate the performance of our TinySAM for segmenting an object from several points as the prompts. We use the same points selection metric as previous work [21, 17], which calculates the distance transform of false positive and false negative masks, and then sample points at a maximal value. We calculate the mIoU of each dataset to evaluate the performance of different models.
@@
Table 2: Comparison of original points grid strategy and our hierarchical strategy. Evaluation on the first 100 images of COCO val2017 set.
Figure 5: Visualization results of COCO validation dataset for zero-shot instance segmentation. The green box marks the box prompt. Our TinySAM captures more clear and smooth boundaries especially for hard targets of small size or similar texture feature.
@@
We choose a subset of total 23 datasets used in [21] for efficient evaluation, which contains BBBC038v1 [2], DOORS [41], TimberSeg [14] and LVIS [17]. To make fair comparisons, we follow the settings of Table 7 in SAM [21] paper to sample the images and masks, and the first $N$ masks in the corresponding split are used in the evaluation.
@@
The evaluation results are shown in Figure 4. Our TinySAM outperforms MobileSAM [50] significantly on LVIS and TimberSeg dataset and obtains similar performance on DOORS dataset. Moreover, our TinySAM achieves better results on BBBC038v1 when fewer number of points is utilized as prompts. We also report the mean IoU of all four datasets, as shown in the right figure in Figure 4. The proposed TinySAM is better than MobileSAM and obtains close performance to that of SAM-B.
Figure 6: Visual comparison of everything mode segmentation results on COCO val2017 set. (a) Points grid strategy result, mIoU 0.5292. (b) Hierarchical strategy result, mIoU 0.6429. Our method avoids the over-fining segmentation of objects in this case.
4.4 Everything Mode Acceleration
@@
We evaluate our proposed hierarchical everything inference strategy on COCO validation set. Latency benchmarks are conducted on a single NVIDIA V100 GPU for everything mode. We sample 100 images with the least img_id from val2017 and conduct everything mode inference on these samples. The threshold values used in the everything mode are all kept the same as default. The results are shown in Table 2. We apply the same threshold and stability score on the same model evaluated with different strategies to make a fair comparison. But they can be different between these models. Our hierarchical strategy achieves comparable results compared with original $32\times 32$ points grid strategy while the cost of inference time is reduced by about $50\%$. Figure 6 shows the visual comparison between points grid strategy and the proposed hierarchical strategy. We can find that the proposed hierarchical strategy avoids the over-finegrained segmentation of objects in some cases which makes up for some losses of IoU. More visual results of the efficient everything mode inference is shown in Figure 7.
Figure 7: Visualization results of hierarchical everything inference with TinySAM model.
4.5 Ablation Studies
@@
In this section, we conduct ablation studies of the proposed method on zero-shot instance segmentation task on COCO validation dataset. The experimental setting is the same as described in Section 4.2.
@@
Impacts of different modules. We first evaluate the effects of different modules, i.e., full-stage knowledge distillation loss, online hard prompt sampling and post quantization, respectively. As shown in Table 3, utilizing our proposed full-stage distillation strategy improve the performance from $40.7\%$ to $41.4\%$. Incorporated with the online hard prompt sampling strategy, our method could obtain $0.5\%$ AP gain. With the hard mask weighting loss, the performance can further increase to $42.3\%$. Using post-training quantization results in $0.9\%$ AP degradation but also greatly reduces the computational cost, which is also discussed in the main experiments.
@@
Point density and threshold for hierarchical strategy. In Table. 5, we conduct ablation study with different settings of point density and high-confidence mask threshold $\tau$. More points and higher threshold $\tau$ lead to more precise results but longer inference time. The point density of 2nd round is more sensitive compared to the 1st one. Considering both accuracy and efficiency, the setting in bold is a good balance and used for other experiments of everything inference mode.
@@
Different bits for quantization. We here explore the influence of different bit width. Table 6 reports the average precision on COCO dataset. From the results, we can conclude that quantization to $8$ bits results in only slight performance drop. We also demonstrate the performance by further reducing the quantization bit width to $6$ bits.
Model AP (%) FLOPS (G)
@@
Table 6: Ablation study for different bit width of quantization for zero-shot instance segmentation on COCO validation dataset.
5 Conclusion
In this paper, we propose a framework to push the envelope for segment anything task and obtain a highly efficient model named TinySAM. We first propose a full-stage knowledge distillation method with hard mask weighting and hard prompt sampling strategy to distill a lightweight student model. We also adapt the post-training quantization to the promptable segmentation task and further reducing the computational cost. Moreover, a hierarchical segmenting everything strategy is proposed to accelerate the everything inference by 2×2$\times$2 × with almost no performance degradation. With all these proposed methods, our TinySAM leads to orders of magnitude computational reduction and push the envelope for efficient segment anything task. Extensive experiments on various zero-shot transfer tasks demonstrate the significantly advantageous performance of our TinySAM against counterparts methods. We hope the proposed TinySAM brings beneficial perspective for designing a highly efficient segment anything model.
Figure 8: Visualization results of LVIS v1 dataset for zero-shot instance segmentation. The green box denotes the box prompt. TinySAM generates more accurate and smooth masks compared with FastSAM and MobileSAM.
@@ 5 Conclusion
We provide more visualization results for supplementary material. Figure 8 shows zero-shot instance segmentation on LVIS v1 [17] dataset. For clear presentation, only detected boxes by VitDet-H [25] with scores higher than $0.8$ are prompted into models. LVIS dataset has more fine-grained labels than COCO dataset [28], on which the proposed TinySAM demonstrates greater advantage considering both accuracy and efficiency.
Figure 9 shows the everything inference results by the proposed TinySAM model with hierarchical everything inference and its counterpart algorithms. TinySAM captures clear boundaries and produce more fine-grained masks, whereas MobileSAM [50] and FastSAM [51] sometimes generate fabricated boundaries and masks. TinySAM shows more close performance to the original SAM [21], while consuming significantly less computation cost.

### v3

@@ Introduction
SAM-H (Kirillov et al. 2023) 2976G 2392 46.6 30.8 51.0 61.7 44.7 32.5 57.6 65.5
SAM-L (Kirillov et al. 2023) 1491G 1146 46.2 30.2 50.1 60.5 43.5 31.1 56.3 65.1
SAM-B (Kirillov et al. 2023) 487G 368.8 43.4 28.5 45.5 53.4 40.8 29.1 52.8 60.7
FastSAM (Zhao et al. 2023) 443G 153.6 37.9 23.9 43.4 50.0 34.5 24.6 46.2 50.8
EfficientSAM-Ti (Xiong et al. 2024) 106G 81.0 42.3 26.7 46.2 57.4 39.9 28.9 51.0 59.9
SlimSAM-77 (Chen et al. 2024) 51.7G 110 41.3 25.7 44.9 57.4 38.3 26.7 49.7 59.0
MobileSAM (Zhang et al. 2023) 42.0G 38.4 41.0 24.4 44.5 58.6 37.0 24.7 47.8 59.1
@@
Figure 4: Results of zero-shot points valid mask evaluation. X-axis represents the number of prompts points and Y-axis represents the mIoU across all masks. The proposed TinySAM outperforms MobileSAM and achieves results close to SAM-B.
Zero-shot Points Valid Mask Evaluation
@@
In this section, we evaluate the performance of our TinySAM for segmenting an object from several points as the prompts. We use the same points selection metric as previous work (Kirillov et al. 2023; Gupta, Dollar, and Girshick 2019), which calculates the distance transform of false positive and false negative masks, and then sample points at a maximal value. We calculate the mIoU of each dataset to evaluate the performance of different models.
@@
Table 2: Comparison of original point grid strategy and our hierarchical strategy. Evaluation on the first 100 images of COCO val2017 set.
@@
We choose a subset of total 23 datasets used in (Kirillov et al. 2023) for efficient evaluation, which contains BBBC038v1 (Caicedo et al. 2019), DOORS (Pugliatti and Topputo 2022), TimberSeg (Fortin et al. 2022) and LVIS (Gupta, Dollar, and Girshick 2019). To make fair comparison, we follow the settings of SAM (Kirillov et al. 2023) to sample the images and masks, and the first $N$ masks in the corresponding split are used in the evaluation.
@@
The evaluation results are shown in Figure 4. Our TinySAM outperforms MobileSAM (Zhang et al. 2023) significantly on LVIS and TimberSeg dataset and obtains similar performance on DOORS dataset. Moreover, TinySAM achieves better results on BBBC038v1 when fewer points are utilized as prompts. We also report the mean IoU of all four datasets, as shown in the right of Figure 4. The proposed TinySAM achieves higher mIoU than MobileSAM and obtains close performance to that of SAM-B.
Figure 5: Visualization for the process hierarchical everything strategy. (a) shows the intermediate result of high-confidence regions after 1st sparse prompt points with white mask and remained 2nd dense prompt points with green stars. (b) shows the final segmentation result and the small objects can be accurately segmented.
Everything Mode Acceleration
@@
We evaluate our proposed hierarchical everything inference strategy on COCO validation set. Latency benchmarks are conducted on a single NVIDIA V100 GPU for everything mode. We sample 100 images with the least img_id from val2017 and conduct everything mode inference on these images. The threshold values used in the everything mode are all kept the same as default. The results are shown in Table 2. We apply the same threshold and stability score on the same model evaluated with different strategies to make a fair comparison, but they can be different between these models. Our hierarchical strategy achieves comparable results compared with original $32\times 32$ points grid strategy while the cost of inference time is reduced by about $50\%$. Figure 5 shows the intermediate visual results of the hierarchical strategy. We can see that the 1st round of sparse inference has segmented and removed the large objects, the remained points focus more on the small objects. This self-adaptive hierarchical strategy efficiently reduces the computation redundancy and maintains the high accuracy. More visual results are shown in the appendix.
Ablation Studies
@@
In this section, we conduct ablation studies of the proposed method on zero-shot instance segmentation task on COCO validation dataset. The experimental setting is the same as described in zero-shot instance segmentation.
@@
Impacts of different modules. We first evaluate the effects of different modules, i.e., full-stage knowledge distillation loss, hard prompt sampling, hard mask weighting and post quantization, respectively. As shown in Table 3, utilizing our proposed full-stage distillation strategy improve the performance from $40.7\%$ to $41.4\%$. Incorporated with the online hard prompt sampling strategy, our method could obtain $0.5\%$ AP gain. With the hard mask weighting loss, the performance can further increase to $42.3\%$. Using post-training quantization results in $0.9\%$ AP degradation but greatly reduces the computational cost.
@@
Model AP (%) MACs (G)
@@
Table 6: Ablation study for different bit width of quantization for zero-shot instance segmentation on COCO val set.
Point density and threshold for hierarchical strategy. In Table 5, we conduct ablation study with different settings of point density and high-confidence mask threshold $\tau$. More points and higher threshold $\tau$ lead to more precise results but longer inference time. The point density of 2nd round is more sensitive compared to the 1st one. Considering both accuracy and efficiency, the setting in bold is a good balance and used for other experiments of everything inference.
Different bits for quantization. We here explore the influence of different bit width. Table 6 reports the average precision on COCO dataset. From the results, we can conclude that quantization to $8$-bit results in only slight performance drop. We also demonstrate the performance by further reducing the quantization bit width to $6$.
Conclusion
In this paper, we propose a framework to push the envelope for segment anything task and obtain a highly efficient model named TinySAM. We first propose a full-stage knowledge distillation method with hard mask weighting and hard prompt sampling strategy to distill a lightweight student model. We also adapt the post-training quantization to the prompt-based segmentation task and further reducing the computational cost. Moreover, a hierarchical segmenting everything strategy is proposed to accelerate the everything inference by 2×2$\times$2 × with almost no performance degradation. With all these proposed methods, our TinySAM leads to orders of magnitude computational reduction and push the envelope for efficient segment anything task. Extensive experiments on various zero-shot transfer tasks demonstrate the significantly advantageous performance of our TinySAM against counterpart methods. We hope the proposed TinySAM brings beneficial perspective for designing a highly efficient segment anything model.
Figure 6: Visualization results of COCO validation dataset (upper 3 rows) and LVIS v1 dataset (lower 3 rows) for zero-shot instance segmentation. The green box marks the box prompt from the ViTDet-H detector. TinySAM captures more clear and smooth boundaries especially for hard targets of small size or similar texture feature.
@@ Conclusion
We provide more visualization results for the appendix. Figure 6 shows zero-shot instance segmentation on COCO dataset (Lin et al. 2014) and LVIS v1 (Gupta, Dollar, and Girshick 2019) dataset, respectively. For clear presentation, only detected boxes by ViTDet-H (Li et al. 2022c) with higher confidence scores than $0.8$ are prompted into models and visualized on the figure. LVIS v1 dataset has more fine-grained mask labels than COCO dataset (Lin et al. 2014), on which the proposed TinySAM demonstrates greater advantage in terms of both accuracy and efficiency.
Figure 7 shows the everything inference results by the proposed TinySAM model with hierarchical everything inference and its counterpart algorithms. TinySAM captures clear boundaries and produces more fine-grained masks, whereas MobileSAM (Zhang et al. 2023) and FastSAM (Zhao et al. 2023) sometimes generate fabricated boundaries and masks. TinySAM shows more close performance to the original SAM (Kirillov et al. 2023), while consuming significantly less computation cost.